. If there are 16 total bottles of diet soda, 8 total bottles of regular soda, and 4 total bottles of water, what is the probability of each of the following:
(i) Choosing a bottle of diet soda when a bottle is chosen at random
(ii) Choosing a bottle of regular soda when a bottle is chosen at random
(iii) Choosing a bottle of water when a bottle is chosen at random
**Answers I came up w/ are below not sure if they are correct**
i) 2/3
ii) 1/3
iii) 1/6

Answer by ewatrrr(24785)

 
Hi,
28 bottles in ALL,of which: 16 are diet soda, 8 are regular soda, and 4 are water.
Chosing 1 bottle..
P(diet soda) = 16/28
P(reg soda) = 8/28
P(water) = 4/28
reduce the fractions and You are done.